# A Concurrent Generalized Kropina Change

### v1

@@ 3 Almost rationally of a $\phi$-Kropina change
1- A Riemannian metric is a quadratic Finsler metric; therefore, it is classified as a rational Finsler metric.
2- The generalized Kropina metric, which can be viewed as a generalized Kropina change of a Riemannian metric, is listed as an almost rational Finsler metric [11].
3-The $m$-th root metric falls into the category as an almost rational Finsler metric [11]. Moreover, the generalized Kropina change of an $m$-th root metric is an almost rational Finsler metric [11].
@@
×$\times\left\{\frac{1}{m}g_{ij}+\frac{(2m+1)}{m+1}\frac{F^{2}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\ell_{i}\,\ell_{j}-\frac{2F}{\Phi}\left\{\phi_{i}\,\ell_{j}+\phi_{j}\,\ell_{i}\right\}\right\}$
@@
×$\times\left\{\frac{1}{m}\zeta\,a_{ij}+\frac{(2m+1)}{m+1}\frac{\zeta\,a_{rs}y^{r}y^{s}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\zeta\frac{a_{ri}y^{r}a_{kj}y^{k}}{a_{ms}y^{m}y^{s}}-\frac{2\zeta\,y^{r}}{\Phi}\left\{\phi_{i}\,a_{rj}+\phi_{j}\,a_{ir}\right\}\right\}$
@@
×$\times\left\{\frac{1}{m}a_{ij}+\frac{(2m+1)}{m+1}\frac{a_{rs}y^{r}y^{s}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\frac{a_{ri}y^{r}a_{kj}y^{k}}{a_{ms}y^{m}y^{s}}-\frac{2y^{r}}{\Phi}\left\{\phi_{i}\,a_{rj}+\phi_{j}\,a_{ir}\right\}\right\}.$
@@ Conclusion
We have investigate what we call the $\phi$-concurrent generalized Kropina change $\widehat{F}$ of an arbitrary Finsler metric $F$ and find the associated Finslerian geometric objects of $\widehat{F}$ in terms of those of $F$ in Proposition 2.5 and Theorems 2.8, 2.10. Also, we give a necessary and sufficient condition to make $\widehat{F}$ Finsler metric in Theorem 2.6. This enable us to prove that:
@@
– If the $(1,1)$-tensor $\mathbb{F}$ defined in (2.22) vanishes identically, the the Barthel connection, horizontal and vertical projectors, the curvature of Barthel connection and Berwald connections are invariant. Therefore, the vector filed $\phi$ becomes concurrent with respect to $\mathbb{F}$ (see, Corollary 2.11).
– As a coordinate-study application (§3) of our results, we deduce that the $\phi$-concurrent generalized Kropina change preserves the almost rational property of the initial Finsler metric ${F}$ (in Theorem 3.3).
@@
– Finally, we prove that the $\phi$-concurrent generalized Kropina change ($F\longrightarrow\widehat{F}$) preserves the almost rational property of the initial Finsler metric ${F}$ in §3.
Declarations
  • Competing interests: The authors declare no conflict of interest.
  • Availability of data and material: Not applicable.
  • Funding: Not applicable.
  • Authors’ contributions: The authors have made substantive contributions to the article and assume full responsibility for its content. The authors read and approved the final manuscript.

### v2

@@ 3 Almost rationally of a $\phi$-Kropina change
A Riemannian metric is a quadratic Finsler metric; therefore, it is classified as a rational Finsler metric.
The generalized Kropina metric, which can be viewed as a generalized Kropina change of a Riemannian metric, is listed as an almost rational Finsler metric, see [11].
The $m$-th root metric falls into the category of almost rational Finsler metrics [11]. Moreover, the generalized Kropina change of an $m$-th root metric is an almost rational Finsler metric [11].
@@
$\times\left\{\frac{1}{m}g_{ij}+\frac{(2m+1)}{m+1}\frac{F^{2}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\ell_{i}\,\ell_{j}-\frac{2F}{\Phi}\left\{\phi_{i}\,\ell_{j}+\phi_{j}\,\ell_{i}\right\}\right\}$
@@
$\times\left\{\frac{1}{m}\zeta\,a_{ij}+\frac{(2m+1)}{m+1}\frac{\zeta\,a_{rs}y^{r}y^{s}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\zeta\frac{a_{ri}y^{r}a_{kj}y^{k}}{a_{ms}y^{m}y^{s}}-\frac{2\zeta\,y^{r}}{\Phi}\left\{\phi_{i}\,a_{rj}+\phi_{j}\,a_{ir}\right\}\right\}$
@@
$\times\left\{\frac{1}{m}a_{ij}+\frac{(2m+1)}{m+1}\frac{a_{rs}y^{r}y^{s}}{\Phi^{2}}\,\phi_{i}\,\phi_{j}+2\frac{a_{ri}y^{r}a_{kj}y^{k}}{a_{ms}y^{m}y^{s}}-\frac{2y^{r}}{\Phi}\left\{\phi_{i}\,a_{rj}+\phi_{j}\,a_{ir}\right\}\right\}.$
@@ Conclusion
We have investigate what we call the $\phi$-concurrent generalized Kropina change $\widehat{F}$ of an arbitrary Finsler metric $F$ and find the associated Finslerian geometric objects of $\widehat{F}$ in terms of those of $F$ in Proposition 2.6 and Theorems 2.9, 2.11. Also, we give a necessary and sufficient condition to make $\widehat{F}$ Finsler metric (in Theorem 2.7). This enable us to prove that:
@@
If the $(1,1)$-tensor $\mathbb{F}$ defined in (2.22) vanishes identically, the the Barthel connection, horizontal and vertical projectors, the curvature of Barthel connection and Berwald connections are invariant. Therefore, the vector filed $\overline{\phi}$ becomes concurrent with respect to $\widehat{F}$ (see, Corollary 2.12).
As a coordinate-study application (§3) of our results, we deduce that the $\phi$-concurrent generalized Kropina change preserves the almost rational property of the initial Finsler metric F{F} (in Theorem 3.4).
@@
Finally, we prove that the $\phi$-concurrent generalized Kropina change ($F\longrightarrow\widehat{F}$) preserves the almost rational property of the initial Finsler metric F{F} in §3.